# Turn Every Application into an Agent: Towards Efficient Human-Agent-Computer Interaction with API-First LLM-Based Agents

### v1

@@ Abstract.
Multimodal large language models (MLLMs) have enabled LLM-based agents to directly interact with application user interfaces (UIs), enhancing agents’ performance in complex tasks. However, these agents often suffer from high latency and low reliability due to the extensive sequential UI interactions. To address this issue, we propose AXIS, a novel LLM-based agents framework prioritize actions through application programming interfaces (APIs) over UI actions. This framework also facilitates the creation and expansion of APIs through automated exploration of applications. Our experiments on Office Word demonstrate that AXIS reduces task completion time by 65%-70% and cognitive workload by 38%-53%, while maintaining accuracy of 97%-98% compare to humans. Our work contributes to a new human-agent-computer interaction (HACI) framework and a fresh UI design principle for application providers in the era of LLMs. It also explores the possibility of turning every applications into agents, paving the way towards an agent-centric operating system (Agent OS).
1. Introduction
As personal computer, mobile devices, and internet become an indispensable part of every day’s work and life, application industries are under great pressure to rapidly evolve their software applications with more features and functionalities to meet peoples’ growing demand (Ruparelia, 2010; Abrahamsson et al., 2017). Nonetheless, those new applications also demand a much higher investment in both time and cognitive effort from regular users. To learn how to use a new application effectively, Users generally have to firstly spend significant time just to get familiar with the user interface (UI) and corresponding functionalities. And to complete various tasks with the new application efficiently, users need to further invest time and effort to learn how to break complex tasks into steps and use the right UIs and commands to complete each step. While both the application providers and research community are fully aware of this pain point, existing efforts have been focusing on providing detailed tutorials and establishing engaging learning platforms, which could only provide limited support in alleviating users’ cognitive burden (Van Merrienboer and Sweller, 2005; Biswas et al., 2005; Plass et al., 2010; Darejeh et al., 2022).
@@ 1. Introduction
Large language models (LLMs) (Ouyang et al., 2022; Achiam et al., 2023; Dubey et al., 2024) has demonstrated near-human capabilities in reasoning, planning, and collaboration and are highly promising in completing complex tasks (Huang and Chang, 2022; Wei et al., 2022; Mandi et al., 2024). Since then, researchers has been exploring how LLMs can be utilized to reduce users’ cognitive burden in learning and operating software applications. In particular, multimodal large language models (MLLMs) (Yin et al., 2023; Durante et al., 2024; Zhang et al., 2024c) expand the usage scenarios of LLMs to various tasks that require vision capability (Wu et al., 2023; Zheng et al., 2024). Recent works (Zhang et al., 2023b; Wang et al., 2024c; Zhang et al., 2024a; Zheng et al., 2024) utilizes MLLMs to design LLM-based UI agents capable of serving as users’ delegates, translating user requests expressed in natural language, and directly interacting with the UI of software applications to fulfill users’ needs. With the help of LLM-based UI agents, users could simply ask the application to complete tasks without a deep understanding of application’s UIs and functionalities, which significantly reduces users’ cognitive load of learning new applications.
@@
However, just like the transition from steam-powered to electric-powered industry took much more than replacing central steam engines with electric motors in the factories, simply building LLM-based agent upon the UIs of applications cannot magically deliver a satisfied and worry-free user experience. In particular, today’s application UIs are designed for human-computer interaction (HCI) (Lewis, 1998; Bradshaw et al., 2017), which often involves multiple UI interactions for completing a single task. For instance, inserting a 22 table in an Office Word document requires a sequence of UI interactions: “Insert Table 22 Table”. Although the HCI-based design suits the habits of humans, training LLM-based UI agents to emulate such interactions would generate quite a few challenges that are difficult to overcome.
@@
The first challenge for the LLM-based UI agent is the high latency and long response time. Each individual UI interaction step requires one LLM call to reason which UI to interact with. A task involves multiple UI interaction steps can thus incur considerable time and monetary cost. The LLM call latency is also positively correlated with the number of processed tokens (Levy et al., 2024; Wang et al., 2024a; Egiazarian et al., 2024). To ensure that the LLM can return high quality outputs, the LLM-based UI agent must pass large volume of UI information to precisely describe the current state, which also increases the latency in each call. The second challenge lies in the reliability domain. Studies have shown that LLMs are prone to hallucinations in generating responses (Bang et al., 2023; Dhuliawala et al., 2023; Zhang et al., 2023a; Guan et al., 2024). During the long sequential calls with LLM-based UI agents, the chance of taking a wrong UI control or hallucinating a non-existing UI for interaction increases with each reasoning step. As LLM-based UI agents often pass the previous interaction history as additional context when reasoning on current UI interaction step (Zhang et al., 2024a), hallucinations in earlier steps could also increase the chance of hallucination in later steps. Thus, when a long chain of UI interactions is required, the UI agents are more likely to suffer from compounding errors and encounter task failure (Chen et al., 2024; Zhao et al., 2024). Lastly, the LLM-based UI agent also faces the challenge of UI generalization. While recent works had made advancements in UI grounding (Cheng et al., 2024; Rawles et al., 2024b; Bai et al., 2023), how the LLM-based UI agents handle interactions with applications whose UIs are not included in the pretraining stage of LLMs remain a critical obstacle without good solutions.
We believe that a new human-agent-computer interaction (HACI) paradigm is needed to address the challenges faced by LLM-based UI agents. In HACI paradigm, API-first LLM-based agents will replace UI agents to prioritize API calls over unnecessary multi-step UI interactions in task completion. Regular UI interactions are only called when the related APIs are unavailable. Compared to the UI agents, API-first agents require less tokens and can obtain more accurate code-formated responses from LLMs. For instance, in the task of inserting a 22 table in Word document, the API call only requires one line of code, i.e., doc.Tables.Add(NumRows=2,NumColumns=2) to complete the task.
In this paper, we propose AXIS: Agent eXploring API for Skill integration, a self-exploration LLM-based framework capable of automatically exploring existing applications, learning insights from the support documents and action trajectories, and constructing new APIs111The new APIs are also referred as “Skills” in Section 3, and we use the term “API” loosely here to differentiate from the UI. based on the existing APIs to empower API-first LLM-based agents with low latency and high reliability. Based on our experiment on Office Word (Microsoft365, 2024a) tasks, AXIS can significantly improve the task completion rate and reduce the cognitive loads of users. Moreover, AXIS provides a practical approach for the application providers to turn an application into agent by simply wrapping up the application with an API set and adopting a simpler UI design suitable for HACI. This application-as-an-agent paradigm also paves the way towards the agent operating system (Agent OS) (Zhang et al., 2024b; Mei et al., 2024; Wu et al., 2024), where the users only need to communicate their intention in natural language and the Agent OS then automatically forms actionable plans, distributes the sub-tasks to relevant applications, and oversees the execution and completion of the tasks with minimal interventions from users.
@@
We propose an HACI paradigm along with the implementation framework named AXIS for creating API-first LLM-based agents capable of exploring the application and its available APIs and constructing new APIs. This new paradigm provides a practical approach to turn every application into an agent and paves the way of developing a real Agent OS.
We address the cognitive load and learning effort challenge by reducing unnecessary multi-step UI interactions and simplifying task completion through API calls.
@@
We conduct performance evaluations and extensive user study to fully examine the efficiency and reliability of AXIS.
2. Related Work
2.1. LLM-based UI Agent
LLM-based agents are designed to utilize the advanced context understanding and reasoning skills of LLMs to interact with and manipulate environments with human-like cognitive abilities († et al.(2022), FAIR; Xi et al., 2023; Liu et al., 2023; Wang et al., 2024b). The advent of MLLM (Yin et al., 2023; Durante et al., 2024; Zhang et al., 2024c), including GPT-4o (OpenAI, 2024a) and Gemini (Team et al., 2023), further broadens the scope of LLMs in practical applications with the capability of processing multi-modal inputs, including text and images. Supported by the vision understanding capability of MLLM, new LLM-based UI agents can acquire crucial abilities of navigating and controlling UIs in software applications for completing complex tasks. As such, study on LLM-based UI agents has emerged as a hot area for developing intelligent assistants that can automatically interact with applications following users’ commands. In the mobile platform, methods such as MM-Navigator (Yan et al., 2023), AppAgent (Zhang et al., 2023b), and MobileAgent (Wang et al., 2024c) leverage GPT-4V (OpenAI, 2024b) to operate smartphone applications through human-like interactions (tapping and swiping) without requiring back-end access. More broadly, UFO (Zhang et al., 2024a), SeeAct (Zheng et al., 2024) and Cradle (Tan et al., 2024) support the navigation and operation of UIs in Windows OS applications, websites and games respectively following commands in natural language. Other notable examples include CogAgent (Hong et al., 2024) and SeeClick (Cheng et al., 2024), which focus on the training and finetuning of visual language models for UI understanding and navigation in downstream mobile and desktop tasks. While these LLM-based UI agents are trained to complete tasks in a human-like manner, UIs in existing applications were originally designed for HCIs rather than for agent-computer interactions. Consequently, emulating UI-based interactions directly can result in unnecessary time costs, especially in complex tasks that require numerous or repeated UI interaction steps, such as changing multiple titles to the same format. In contrast, application providers usually offer APIs that can accomplish such tasks with a single API call that eliminate the need of multi-step UI interactions. To overcome the limitation of existing UI agents, reduce unnecessary UI interactions, and lower humans’ learning curve for applications, we will study how the application APIs can be leveraged for building LLM-based agents and explore the new design principles of UIs in the era of LLMs.
2.2. Agent Operating System
The LLM-based agents discussed in Section 2.1 are usually designed to work within a narrow environment such as a specific application or web page which limits their applicability in general computer tasks that often involves cross-application collaboration. For example, a simple task “Create a report with docs under the Presentation folder and send it to Jack.” requires multiple steps with multiple applications to complete: read all docs, summarize key contexts, compose a report with a word processing application like Microsoft Word, and draft an email with attached report using an email client like Outlook, and sent it to the recipient Jack. To support the completion of complex tasks with minimal human interventions, emerging works have explored the possibility of developing an operating system (OS) fully supported by LLMs. Works like AIOS (Mei et al., 2024) and OS-Copilot (Wu et al., 2024) propose an OS-level agent that can effectively interact with the OS and a vast number of third-party applications in completing complex tasks. OSWorld (Xie et al., 2024) and AndroidWorld (Rawles et al., 2024a) also provide benchmarks for evaluating the performance of multimodal agents with diverse tasks and cross-application workflows across various OS. In the industry, commercial Agent OS such as Apple Intelligence (Apple, 2024), Copilot+PC (Microsoft, 2024), HarmonyOS (Huawei, 2024), and MagicOS (Honor, 2024), are evolving to be more accessible and productive for customers with the potential of leading a new era of HCI. A common approach adopted by existing Agent OS is to divide complex into sub-tasks and assign them to individual applications. However, for each sub-task, the LLM-based agents still rely on human-like interactions such as UI clicking and swiping for completion, which can be inefficient compared to API calls. Moreover, when the LLM-based UI is processing the task, the control is taken over from the user by the LLM-based agent.
2.3. UI design in LLM era
@@
UI design is an essential part of HCI and requires highly specialized expertises along with iterative rounds of feedback and revision (Stone et al., 2005). With LLMs, UI design can be further empowered with automated procedures of design, feedback and evaluation. Duan et al. (2024) use LLM-generated feedback to automaticly evcuate UI mockups. Similiarly, SimUser (Xiang et al., 2024) leverage LLMs to simulate users with different characters to generate feedback on usability and provide insights in UI design. MUD (Feng et al., 2024) utilize LLMs to mimic human-like exploration to mine UI data from applications and employs noise filtering to improve quality of UI data. Still, existing UI designs largely follow the traditional human-machine interaction paradigm rather than the human-agent-computer interaction (HACI) paradigm that could became the central design principle in Agent OS. In this paper, we will leverage AXIS framework to explore applications, identify the essential UIs, and examine which part of UIs can be replaced by APIs callable by LLM-based agents. We hope that our exploration would generate insights on how to design UIs that are more effective under the HACI paradigm.
3. Design of AXIS
We develope AXIS as a framework that can automatically explore within existing applications, learn insights from exploration trajectories, and consolidate available insights and learned knowledge into actionable “skills”. Skill is a high-level representation of UI- and API-based actions with a priority in API actions222If the skill can be represented with UI or API actions, the skill is represented in API-only actions., which is generated with AXIS exploration within the application. Illustrated by Figure 2, AXIS system consists of three major modules: the app environment, the skills, and the workflows. AXIS employs numerous LLM-based agents to explore the app environment, usually a set of applications running on the OS, through a unified interface to obtain the state of the environment and to interact with it. The knowledge learnt during this process will be consolidated into skills containing structured code segments capable of accomplishing various tasks within the environment. Specific execution and validation methods will also be designed to improve the performance of those skills. Finally, we establish two workflows: the explorer workflow and the follower workflow to facilitate the learning of skills from the environment.
In the subsequent sections, we will layout the details of all the 3 modules and discuss how they work together to explore and discover valuable skills from the applications.
3.1. Environment of Application
@@
In the context of AXIS, the environment of applications refers to the collection of interactive entities within the exploration scope of the agents. In this paper, those entities mainly consist of a set of applications running on the Windows operating system. Applications in the environment often share common elements, such as controls (Zhang et al., 2024a) and XML elements obtained after unpacking.
Agents in AXIS not only observe the state of the environment, but also actively interact with the environment. To facilitate the observation and interaction between agents and the environment, we have designed two general interfaces: state() and step(). state() interface returns the state of the environment, which includes detailed information on the current elements of the entities within the environment. The environment state encompasses key UI information including the position of controls, the type of controls, and whether a control is selected. For applications that can be unpacked, the unpacked XML content is also included as a part of the state. The control types in the application environment of AXIS are consistent with those in (Zhang et al., 2024a). step() interface incorporates a skill executor that allows agents to perform operations within the environment by executing skills. Upon completion, this interface also returns the results of these operations.
3.2. Skills in Application
A skill contains skill code, description, and usage example, and is designed to accomplish one specified task within the environment.
3.2.1. skill structure
Skill Code: a piece of code structured to be compatible with the executor described in the following section. Skill code includes a uniform set of parameters and adheres to the standard PEP 257 documentation. The initial set of skills is generated by the restructure of the fundamental APIs from the application provider. Based on these initial skills, AXIS can explore and develop additional new skills. It is worth noting that while AXIS prioritizes API-based skills, it is still designed for general purpose and can incoporate both UI-based and API-based skill code.
Description: A description of the functionality of a skill for assisting the LLM in selecting and invoking the appropriate skil in the process of task execution.
Usage Example: One or more code examples including any specific parameters typically associated with the code and the description. These examples can assist the LLM in filling out the parameter fields in the correct format when the skill is invoked.
3.2.2. skill executor
As discussed in Section 3.1, our application environment incorporates a step() interface to facilitate the interaction between agents and the environment. This interface also hosts the skill executor responsible for executing the skill generated or selected by the agents. The skill executor keeps caching of application documents and simultaneously supports multiple functionalities including locating application controls, invoking methods on those controls, and calling application APIs (independent of controls), to enable the UI actions and API actions in the same time and serve as an efficient foundation for skill-driven operations.
3.2.3. skill types
@@ 3.2. Skills in Application
Following a versatile design principle, the skills in AXIS can be categorized into four types based on the composition of their code fragments: Atomic UI Skill, Atomic API Skill, Composite UI Skill, Composite API Skill, and API-UI Hybrid Skill.
| Type | Description | Example | Feature coverage |
| Composite API skill | Composed of multiple atomic API actions or composite API actions. This type of skills often represents a higher-level combination of functions. | insert_header_footer | insert header and footer with specified contents by API, which is equal to sequential UI actions ”Insert-¿Header-¿footer edit-¿Footer-¿footer edit”. |
3.2.4. Skill Hierarchy
We define “skill hierarchy” as the number of skill components that make up a skill. A single basic skill thus has a skill hierarchy of 1. The skill hierarchy of insert_header_footer skill mentioned in table 1 is 2.
3.3. Workflows of Skill Exploration
As shown in Figure 2, the skill exploration in AXIS is guided by two driving mechanisms: follower-driven skill exploration and explorer-driven skill exploration.
3.3.1. Follower-driven skill exploration
Follower-driven skill exploration refers to the process of exploring skills from the applications’ help documents, which is primarily accomplished through the collaboration of the following agents:
FollowerAgent. FollowerAgent utilizes a skill library composed of a set of primitive actions based on the step-to-step instructions provided by the help document. At each step, it selects the most appropriate action according to the current state of the environment.
MonitorAgent. MonitorAgent monitors every action taken by the FollowerAgent along with the impacts on the environment. It also tracks the entire trajectory of the FollowerAgent. When deemed appropriate, it places a breakpoint and summarizes the observed trajectory into a complete skill including the functional summary of the skill and its logic. For initial skills composed of basic UI actions, the summarized logic is also often closely tied to the UI.
SkillGeneratorAgent. Based on the summaries and logical descriptions provided by the MonitorAgent, SkillGeneratorAgent generates the code, functional descriptions, and use cases for the skill in accordance with the specifications outlined in Section 3.2. The generated skill code is highly correlated with the logical descriptions generated by the MonitorAgent, and therefore often involves the stacking of basic UI actions.
APITranslatorAgent. Based on the skill code generated by the SkillGeneratorAgent, APITranslatorAgent queries the relevant API documentation to translate the UI-based actions in the skill code into API calls, thus completing the API-ification of the code. Finally, the generated skills are further validated through a skill validation process.
It is worth noting that both the SkillGeneratorAgent and the APITranslatorAgent search for reusable skills from the original skill set during the skill code generation, thereby obtaining skills at different hierarchies. This approach efficiently facilitates the construction of skills.
3.3.2. Explorer-driven skill exploration
Unlike Follower-driven skill exploration, the explorer-driven skill exploration kicks off the exploration process with a different initialization method: the step-to-step instructions are automatically generated rather than extracting from the help document. During each step, explorer proposes the next action based on the current environment information and the history of previously explored steps. The subsequent steps utilize the same agents in the follower-driven mode to generate skills with explored trajectories. In this process, application seed files with varying pre-filled content are often required to obtain different initial environments for discovering a more diverse range of skills.
3.3.3. Skill Validation
To validate the new skills generated from exploration, we have implemented two verification methods: static validation and dynamic validation.

### v2

# AXIS: Efficient Human-Agent-Computer Interaction with API-First LLM-Based Agents

@@ Abstract
Multimodal large language models (MLLMs) have enabled LLM-based agents to directly interact with application user interfaces (UIs), enhancing agents’ performance in complex tasks. However, these agents often suffer from high latency and low reliability due to the extensive sequential UI interactions. To address this issue, we propose AXIS, a novel LLM-based agents framework that prioritize actions through application programming interfaces (APIs) over UI actions. This framework also facilitates the creation and expansion of APIs through automated exploration of applications. Our experiments on Microsoft Word demonstrate that AXIS reduces task completion time by 65%-70% and cognitive workload by 38%-53%, while maintaining accuracy of 97%-98% compared to humans. Our work contributes to a new human-agent-computer interaction (HACI) framework and explores a fresh UI design principle for application providers to turn applications into agents in the era of LLMs, paving the way towards an agent-centric operating system (Agent OS).
AXIS: Efficient Human-Agent-Computer Interaction with
API-First LLM-Based Agents
Junting Lu1*, Zhiyang Zhang2††thanks: Equal Contribution. Work is done during the internship at Microsoft., Fangkai Yang3††thanks: Corresponding author., Jue Zhang3, Lu Wang3, Chao Du3, Qingwei Lin3, Saravan Rajmohan3, Dongmei Zhang3, Qi Zhang3 1Peking University, 2Nanjing University, 3Microsoft aidan.lew.37@stu.pku.edu.cn
1 Introduction
@@
As personal computers and mobile devices become indispensable in daily life, application industries face pressure to rapidly evolve software with new features to meet growing demands Ruparelia (2010); Abrahamsson et al. (2017). However, to use a new application effectively, users must first spend time familiarizing themselves with the user interface (UI) and its functionalities, increasing users’ cognitive burden Van Merrienboer and Sweller (2005); Biswas et al. (2005); Plass et al. (2010); Darejeh et al. (2022). Large language models (LLMs) Ouyang et al. (2022); Achiam et al. (2023); Dubey et al. (2024) have demonstrated near-human capabilities in reasoning, planning, and collaboration and are highly promising in completing complex tasks Huang and Chang (2022); Wei et al. (2022); Mandi et al. (2024). Since then, researchers have been leveraging multimodal large language models (MLLMs) Yin et al. (2023); Durante et al. (2024); Zhang et al. (2024b) to operate software applications with vision capability Wu et al. (2023); Zheng et al. (2024). Recent works Zhang et al. (2023a); Wang et al. (2024a); Zhang et al. (2024a); Zheng et al. (2024) utilize MLLMs to design LLM-based UI agents capable of serving as user delegates, translating user requests expressed in natural language, and directly interacting with the application’s UIs to fulfill users’ needs without a deep understanding of UIs and functionalities.
@@
However, just like the transition from steam-powered to electric-powered industry took much more than replacing central steam engines with electric motors in the factories, simply building LLM-based UI agent cannot magically deliver a satisfying and worry-free user experience. In particular, today’s application UIs are designed for human-computer interaction (HCI) Lewis (1998); Bradshaw et al. (2017), which often involves multiple UI interactions for completing a single task. For instance, inserting a 22 table in an Microsoft Word document requires a sequence of UI interactions: “Insert Table 22 Table”. Although the HCI-based design suits the habits of humans, training LLM-based UI agents to emulate such interactions would pose quite a few challenges that are difficult to overcome.
The first challenge for the LLM-based UI agent is high latency and long response time. Each individual UI interaction step requires one LLM call to reason which UI to interact with. A task involves multiple UI interaction steps can thus incur considerable time and monetary costs. The LLM call latency is also positively correlated with the number of processed tokens Levy et al. (2024); Wang et al. (2024d); Egiazarian et al. (2024). To ensure the LLM returns high-quality outputs, the LLM-based UI agent must pass a large volume of UI information to precisely describe the current state, which also increases latency in each call.
@@
The second challenge lies in the domain of reliability. Studies have shown that LLMs are prone to hallucinations in generating responses and selecting correct UIs when a long chain of UI interactions is required Bang et al. (2023); Dhuliawala et al. (2023); Zhang et al. (2023b); Guan et al. (2024). During long sequential calls, the chance of selecting a wrong UI control or hallucinating a non-existent UI for interaction increases with each reasoning step Zhang et al. (2024a), resulting in compounding errors and task failure Chen et al. (2024); Zhao et al. (2024). Lastly, the LLM-based UI agent also faces the challenge of UI generalization. While recent works had made advancements in UI grounding Cheng et al. (2024); Rawles et al. (2024b); Bai et al. (2023), how the LLM-based UI agents handle interactions with applications whose UIs are not included in the pretraining stage of LLMs remains a critical obstacle without good solutions.
To address these challenges faced by LLM-based UI agents, we highlight the necessity of application programming interfaces (APIs) within the new human-agent-computer interaction (HACI) paradigm. As an intermediate layer between human and computer, the LLM-based agent should understand user requests in natural language and operate computers/applications by prioritizing APIs over human-like UI interactions. Inspired by this, we propose AXIS: Agent eXploring API for Skill integration, a self-exploration LLM-based framework capable of automatically exploring existing applications, learning insights from the support documents and action trajectories, and constructing new APIs111The new APIs are also referred as “Skills” in Section 4, and we use the term “API” loosely here to differentiate from the UI. based on the existing ones. AXIS helps build API-first LLM-based agents that replace UI agents to prioritize API calls over unnecessary multistep UI interactions in task completion. Regular UI interactions are only called when the related APIs are unavailable. Figure 1 shows a task completion with the UI agent and API-first agent. Compared to the UI agents, API-first agents require fewer tokens and can obtain more accurate, code-formatted responses from LLMs with low latency and high reliability.
@@
We propose a novel HACI paradigm along with an implementation framework, AXIS, which enables API-first LLM-based agents to explore an application’s available APIs and construct new ones as needed. This approach facilitates the straightforward transformation of any application into an autonomous agent by wrapping it with an API set and adopting a simplified UI design. This paradigm not only addresses practical challenges faced by application providers but also paves the way toward a full-fledged Agent Operating System (Agent OS) Zhang et al. (2024c); Mei et al. (2024); Wu et al. (2024).
We mitigate cognitive load and reduce the learning effort required for complex interactions by replacing multi-step UI operations with efficient API calls. Our experiments on Microsoft Word tasks Microsoft365 (2024a) show that AXIS significantly improves task completion rates while alleviating the cognitive burden on users.
@@
We conduct comprehensive performance evaluations and an extensive user study to validate the efficiency, reliability, and practical applicability of AXIS in real-world scenarios.
2 Related Work
2.1 LLM-based UI Agent
LLM-based agents are designed to utilize the advanced context understanding and reasoning skills of LLMs to interact with and manipulate environments with human-like cognitive abilities † et al.(2022)(FAIR)†, Bakhtin, Brown, Dinan, Farina, Flaherty, Fried, Goff, Gray, Hu et al. (FAIR); Xi et al. (2023); Liu et al. (2023); Wang et al. (2024b). The advent of MLLMs Yin et al. (2023); Durante et al. (2024); Zhang et al. (2024b), including GPT-4o OpenAI (2024) and Gemini Team et al. (2023), expands the research landscape for LLM-based UI agents. LLM-based UI agents have been applied to multiple areas such as mobile platforms Yan et al. (2023); Zhang et al. (2023a); Wang et al. (2024a), Web Song et al. (2024) and OS Wang et al. (2024c); Zhang et al. (2024a); Zheng et al. (2024); Tan et al. (2024); Hong et al. (2024); Cheng et al. (2024). LLM-based UI agents mimic human interactions, but existing UIs are designed for human-computer, not agent-computer, interactions, leading to inefficiencies in repetitive tasks. APIs offer a more efficient alternative by reducing unnecessary UI steps. We explore leveraging APIs for LLM-based agents and propose new UI design principles for the LLM era.
2.2 Agent Operating System
To support the completion of complex tasks with minimal human interventions, emerging works have explored the possibility of developing an agent operating system (Agent OS) fully supported by LLMs Mei et al. (2024); Wu et al. (2024); Zhang et al. (2024c); Xie et al. (2024); Rawles et al. (2024a). In the industry, commercial Agent OSes Apple (2024); Microsoft (2024); Huawei (2024); Honor (2024) are evolving to become more accessible and productive for customers with the potential of leading a new era of HCI. Existing Agent OSes typically divide complex tasks into sub-tasks and assign them to applications. However, LLM-based agents still rely on human-like UI interactions, such as clicking and swiping, which are less efficient than API calls. Additionally, when processing a task, the LLM-based UI takes control away from the user.
2.3 UI design in LLM era
@@
UI design is an essential part of HCI and requires highly specialized expertise along with iterative rounds of feedback and revision Stone et al. (2005). With LLMs, UI design can be further empowered with automated procedures of design, feedback and evaluation. Duan et al. (2024) use LLM-generated feedback to automatically evaluate UI mockups. Similiarly, SimUser Xiang et al. (2024) leverages LLMs to simulate users with different characteristics to generate feedback on usability and provide insights into UI design. MUD Feng et al. (2024) utilize LLMs to mimic human-like exploration to mine UI data from applications and employs noise filtering to improve quality of UI data. Existing UI designs follow the traditional HCI paradigm rather than the HACI paradigm central to Agent OS. Using the AXIS framework, we explore applications, identify essential UIs, and determine which UI components can be replaced by API calls for LLM-based agents.
3 Preliminary
@@
Environment. In the context of AXIS, the environment refers to the collection of interactive entities within the exploration scope of the agents. In our case, these entities primarily consist of applications running on the Windows operating system, with Microsoft Word being our focus. Applications in the environment often share common elements, such as controls Zhang et al. (2024a) and XML elements obtained after unpacking. To facilitate the observation and interaction between agents and the environment, we have designed two general interfaces: state() to return the environment state and step() to execute agent actions, respectively. Details of interfaces are in Appendix A.
Skills. A skill in AXIS is a structured unit designed to accomplish a specific task within the environment. It is a high-level representation of UI- and API-based actions, with priority given to API actions222If the skill can be represented with UI or API actions, the skill is represented in API-only actions.. Following the design of tool usage and function call Cai et al. (2023); Wang et al. (2023), each skill consists of three components: description, skill code, and usage example. Appendix B.1 shows details of each component.
@@
Skill Types and Hierarchy. Following a versatile design principle, the skills in AXIS can be categorized into five types based on the composition of their code fragments: Atomic UI Skill, Atomic API Skill, Composite UI Skill, Composite API Skill, and API-UI Hybrid Skill, the details are shown in Appendix B.2 Table 6. Additionally, we define “skill hierarchy” as the number of skills contained. A single atomic skill thus has a skill hierarchy of 1. Based on their code composition and hierarchy, skills in AXIS can be nested. For example, skill A can call skill B, which in turn calls skill C, forming a hierarchy of depth 3.
4 Design of AXIS
We develop AXIS as a framework that can automatically explore within existing application environments, learn insights from exploration trajectories, and consolidate available insights and learned knowledge into actionable “skills”. Illustrated by Figure 2, the AXIS system is composed of three crucial stages: trajectory collection, skill generation, and skill validation. The trajectory collection stage collects interaction trajectories in task completion, and then the skill generation stage generates skills from these trajectories and translates into skill code. The skill validation stage validates the skill code before it is added in the skill library to reduce hallucination and maintain generalizability.
4.1 Stage I: Trajectory Collection
“Experience is the mother of wisdom”. Following the approach in Wang et al. (2023), we let the agent practice acquiring skills in the application environment. In this stage, we designed two practice modes: Follower Mode and Explorer Mode, which are task-oriented collections with and without task descriptions, respectively. The specific prompts refer to Appendix LABEL:sec:traj-prompts.
Follower Mode. The Follower Agent extracts tasks and instructions from the application help documents. The agent processes structured inputs: task requirements, step-by-step instructions, environmental context, and the current skill library. It strictly adheres to instructions while interacting with seed application files, collecting execution trajectories. Action selection employs the ReAct mechanism Yang et al. (2023), integrating action history, step descriptions, and environmental states. Initially limited to basic actions (see Table 7), the follower agent is able to perform more complex tasks with the expansion of the skill library through the other parallel processing stages.
Explorer Mode. Explorer mode differs from the follower mode by leveraging the brainstorming capability of LLMs to generate various step-by-step instructions. The Explorer agent observes environment states and action history, selecting actions from the skill library without following predefined task guides. To enhance state diversity, we implemented:
Initial State Diversity: Random replay of Microsoft Word files up to intermediate steps, providing varied starting points.
Vertical/Horizontal Exploration: A random “dive into” strategy balancing continuity (vertical) and breadth (horizontal) in functionality exploration. “Vertical” means the agent will select UI controls at lower levels of the current control tree for its next action, while “Horizontal” means the agent will tend to select UI controls at the same level or switch to other menus.
Skill Proficiency Levels: Three explorer levels (low/medium/high) guided by Microsoft Office Specialist (MOS) certification curriculum333https://learn.microsoft.com/en-us/credentials/certifications/mos-word-2019/, with system prompts tailored to their app familiarity.
4.2 Stage II: Skill Generation
This stage runs in parallel with Stage I and is responsible for converting the trajectories into structured skills. This process involves three LLM-based agents. The specific prompts for these agents are detailed in Appendix LABEL:sec:skillgen-prompts.
Monitor. The collected trajectories often contain complex and multiple interaction steps, making it challenging to convert them into a single skill without enough basic skills available in the skill library. The monitor examines the skill library, extracts meaningful trajectory segments, and consolidates them into natural language-based skill insights.
Generator. The Generator agent faithfully combines the skill insights with their corresponding trajectory segments and generates executable skill code. As shown in Figure 2, the skill code mainly consists of stacked actions from the trajectories, where the original action parameters are converted into placeholders. Along with the skill code generation, it also produces corresponding skill usage examples and descriptions, with examples shown in Table LABEL:tab:skill-trans. Note that since the skill code is a faithful reproduction of the trajectory, which may contain numerous UI operations, the skill code essentially becomes a combination of UI actions, leaving room for speed optimization.
Translator. As the skill code from the generator contains numerous UI operations, a “UI” to “API” translation is needed. The Translator agent accomplishes this task by connecting to the retrieval-augmented generation (RAG) module Gao et al. (2023), consulting the application’s documentation and the current skill library to convert UI operation code segments to APIs. As shown in Figure 2, it converts UI operations into direct modifications of target content within the application. Table LABEL:tab:skill-trans shows a translation example.
4.3 Stage III: Skill Validation
Due to potential LLM hallucinations and cumulative errors in multi-agent transmission, generated skills may have syntax or functionality issues. To address this, we implement static and dynamic validation, with dynamic validation following static validation.